Providing for Doubtful Accounts 


At the end of the current year, the accounts receivable account has a debit balance of $1,835,000 and sales for the year total $25,690,000 

a. The allowance account before adjustment has a debit balance of $12,500. Bad debt expense is estimated at 1/2 of 1% of sales. 

b. The allowance account before adjustment has a debit balance of $12,500. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $162,000. 

c. The allowance account before adjustment has a credit balance of $26,810. Bad debt expense is estimated at 3/4 of 1% of sales. 

d. The allowance account before adjustment has a credit balance of $26,810. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $171,200. 


Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above. 
Answer #1

Adjusting entry amount

a) Bad debt expense = 25690000*1%*1/2 = 128450

b) Bad debt expense = 162000+12500 = 174500

c) Bad debt expense = 25690000*1%*3/4 = 192675

d) Bad debt expense = 171200-26810 = 144390